OpenStudy (anonymous):

Which line is perpendicular to the line y=2/3x + 3

OpenStudy (mathmate):

Lines that are perpendicular to each other have the product of their slopes equal to -1. For example, y=(3/4)x+5 has a slope of 3/4 y=-(4/3)x-6 has a slope of -4/3 The product of their slopes is (3/4)*(-4/3)=-1, so theses two lines are perpendicular to each other.

OpenStudy (anonymous):

Thank you very much mathmate for your help! I found my answer

OpenStudy (mathmate):

You're welcome! :)